OpenStudy (anonymous):

Prove the trigonometric identity:

OpenStudy (anonymous):

\[(1-\sin^2x)\left( 1+{1 \over \tan^2x} \right)=1\]

OpenStudy (anonymous):

sorry wrong one

OpenStudy (anonymous):

typing it up again

OpenStudy (experimentx):

( 1 + tan^2x) ??

OpenStudy (anonymous):

\[(1-\cos^2x)\left( 1+ {1 \over \tan^2x} \right)=1\]

OpenStudy (anonymous):

The above is the identity I am talking about

OpenStudy (anonymous):

Show your steps please

OpenStudy (campbell_st):

1st brackets are sin^2 2nd brackets are cosec^2 or 1/sin^2 so its\[\sin^2 \times \frac{1}{\sin^2} = 1\]

OpenStudy (anonymous):

Use the Pythagorean identities to help you simplify that:\[\sin^2 x + \cos^2 x = 1\]\[1+\cot^2 x=\csc^2 x\]Start with this substitution: \(\dfrac{1}{\tan^2 x}=\cot^2 x\)

OpenStudy (experimentx):

\[ 1 - cos^2x = \sin^2x\] and \[ 1 + \frac 1 {\tan ^2x} = \frac{\sin ^2x + \cos ^2x }{\sin^2 x} = \frac{1}{\sin^2 x} \]

OpenStudy (anonymous):

How did you do that? Also, make sure you're looking at the new identity, the first one had sin2x instead of cos2x

OpenStudy (anonymous):

Individual steps are appreciated

OpenStudy (anonymous):

I don't get it? @experimentX

OpenStudy (campbell_st):

it all starts with \[\sin^2 + \cos^2 = 1\] subtract cos^2 from both sides \[\sin^1 = 1 - \cos^2\] start again with sin^2 + cos^2 = 1 divide through by sin^2 \[1 + \frac{\cos^2}{\sin^2} = \frac{1}{\sin^2}\] this the the 2nd brackets since \[\frac{1}{\tan^2} = \frac{\cos^2}{\sin^2}\] so you end up with \[\sin^2 \times \frac{1}{\sin^2} = 1\]

OpenStudy (campbell_st):

should be sin^2 = 1 - cos^2

OpenStudy (experimentx):